# Explaining Context Length Scaling and Bounds for Language Models

### v1

@@ 4.5 Optimal Context Length and Training Dataset Size: Synthetic Dataset
To validate the idea that longer context lengths are suitable for larger datasets, we conduct experiments on our proposed synthetic dataset with MLP of fixed size. During our experiments, we change the dataset size and context length used, and obtain validation Binary Cross Entropy loss v.s. Context Length, as shown in the lower figure in Figure 1. (We also plot $y=-\log(1/2)$ as the BCE loss of random guess in Figure 1.)
From Figure 1, we see that models with longer context length need larger datasets to train, and training long-context models on small datasets could lead to worse performance or even failure to improve from the first step (performance equaling to random guess). Moreover, for each curve that represents certain training dataset size used, we observe an optimal horizon beyond which validation loss would increase. This critical point moves right (optimal context length increases) with the increment of training dataset size. This proves Point 3 in Section 4.1.
@@ 5.1 Enlarging Context Length for LMs
Other work modifying architectures has also been proposed to enhance long context modeling or to simplify deployment of long context LLMs. For example, (Tang et al., 2024) proposes a training-free RazorAttention algorithm to largely compress KV cache while maintaining performance unchanged.
Architectures and inference methods have been proposed to lower inference time and memory cost for Language Models, represented by a series of Linear-Transformer or RNN-based methods.(Katharopoulos et al., 2020; Gu & Dao, 2024; Sun et al., 2024). These methods, largely reducing the computational cost and memory usage for long input contexts, has displayed margin ahead of traditional attention based langauge models for long context inference.
Currently a common practice to train very large Language Models supporting long context is to use pretrain the model with shorter contexts, then finetune them with longer contexts, as presented in tech-reports of LLaMa-3(Grattafiori et al., 2024) and DeepSeek-v3(DeepSeek-AI et al., 2024).
@@ 5.2 Irrelevant Long Context hurts performance of LMs
Besides context length scaling with relevant contexts, previous researches have studied how LLMs perform for long irrelevant contexts. As an example, (Levy et al., 2024) studies the performance of current LLMs on an adjusted version of ‘needle in a haystack’ task, where two pieces of key information is embedded into a long text corpora, and a question related to both of them is asked, similar to that presented in Figure 6. The conclusion of these work is that LLMs would perform worse when there is too much irrelevant information.
@@ 5.3 Long Context in another field: Time Series Forecasting
Context length, representing the length of input context, is not unique to Nature Language. For Time Series Forecasting where machine learning plays an important row, there is also work discussing the impact of context length, represented by (Shi et al., 2024). These researches find that there exists an optimal look back horizon, which increases with dataset size. However, time series datasets are relatively small compared to NLP datasets, thus whether this conclusion holds on NLP remains an open problem for this work to study.
@@ 5.4 Related Theories for Scaling Laws
Since the discovery of Scaling Laws for Large Language Models (Kaplan et al., 2020) or even earlier, there have been theoretical work trying to explain why model performance could benefit from more data points and more model parameters. For exmaple, (Sharma & Kaplan, 2022) studies the dataset and model scaling from the data manifold perspective.
Specially for Language Models, there is also previous work proposing all kinds of theoretical models. For example, (Michaud et al., 2024) proposes a feature-quant based theory; (Aghajanyan et al., 2020) views the effect of finetuning from the intrinsic dimension perspective; (Havrilla & Liao, 2024) proposes to understand scaling with intrinsic dimensions.
@@ 6.1 Conclusion
In this work, we discuss the impact of context length to Language Modeling, especially to Bayes Risk and Approximation Loss, from both theoretical and experimental perspective.
In Section 2, we propose assumptions related to the relationship between CE Loss, intrinsic dimension and context length. We derive a linear relation between CE loss and Intrinsic Dimension, and study the impact of context length to intrinsic dimension. We further investigate the relationship between intrinsic dimension, context length and Entropy in Intrinsic Space in Appendix C from an Entropy perspective.
@@ 6.2 Limitations and Future Work
There are limitations of this work. In Section 3 we mainly discuss the case for pretraining with long context, but common practice is to train with shorter contexts and then do long-context-training with longer contexts (DeepSeek-AI et al., 2024). We leave these to future work.
Our theory starting from Intrinsic Dimension only holds with asssumptions in Section 2; and in Appendix C we use the perspective of Entropy and possible states number in Intrinsic Space to (partially) explain our assumptions and measurements w.r.t. Intrinsic Dimension. We hope future work may try to propose even more fundamental theories to explain our Intrinsic Dimension measurements and assumptions.
@@ Appendix A Definition of Cross-Entropy loss used in this work
It is well-known that,
@@
This is because we assume each dimension in the intrinsic space can store $S$ bits of information.
Another intuitive example is: assuming that the vocab is an integer from 00 to $12^{dim(\infty)*S}-$ 1. assuming $P(x_{0}|x_{-\infty:0})=\delta_{x_{0},y}$, that is, the next token given $x_{-\infty:0}$ is sure to be $y$. $y$. For $P_{l}(x_{0}|x_{-\infty:0})$, the first $dim(l)S$ digits of the integer (in binary representation) is known, but the remaining $(dim(\infty)-dim(l))S$ digits are unknown, making a guess in these numbers yield $P_{l}(x_{0}|x_{-\infty:0})=1/2^{S*(dim(\infty)-dim(l))}$. Thus, $D_{KL,x_{0}}(P(x_{0}|x_{-\infty,0}),P_{l}(x_{0}|x_{-\infty,0}))=1*\log 1/(1/2^{S*(dim(\infty)-dim(l))})=S*(dim(\infty)-dim(l))$.
@@ C.1 Bayes Risk from an Entropy-in-Intrinsic-Space perspective
Here we derive similar results in Section 2.2, but not from an Intrinsic-Dimension perspective, rather we derive it from an Entropy perspective.
@@
  • Assumption 2. $\forall l_{1},l_{2}\text{ such that }l_{1}<l_{2}$, $H_{ntp}(l_{1})<H_{ntp}(l_{2})$. This is because longer context contains more information about the next possible token.
@@
  • Assumption 3. ${D_{KL}}_{,ntp}(P,P_{l})\approx H_{ntp}(P)-H_{ntp}(P_{l})+Const$. This means the KL divergence between the Bayes LM of infinite context length and Bayes LM of context length $l$ can be approximately written in the form of entropy difference.
@@
To establish a relationship between Cross Entropy and Intrinsic Space, we run LLaMa-3-8b on a subset of Openwebtext dataset and obtain the feature of the last token as the feature representation, or Intrinsic Space of the approximated Bayes Model, as shown in the left part of Figure 10. We see that model with larger context length tends to have larger relative eigen values in intrinsic space, thus containing more information.
@@
Where $h$ is the ‘plank constant’, meaning that one state corresponds to a unit hyper-volume of $h^{dim}$ in the Intrinsic Space. Different value of $h$ would only add a constant to $H_{IS}$ and would not affect change in Entropy. Thus, we use $\sum_{i}\log rel\_eigval_{i}$ as Entropy in Intrinsic Space. Moreover, experiments in Figure 10 supports that relative eigen values increases at the same ratio as context length increases. This means the number of states in any subspace would scale proportionally w.r.t. number of states in the whole space, that is: $H_{subspace}=H_{IS}*dim(subspace)/dim(IS)$, hence Entropy of subspaces are linear with respect to each other.
Though related, Entropy in Intrinsic Space does not equal to Entropy in next token prediction task. From the probability perspective, let $dec(x)$ be the decoded next token for some point $x$ in the intrinsic space, we have: $H_{IS}=\sum_{x\in IS}-P(x)\log P(x)$, while $H_{ntp}=-\sum_{v\in vocab}P(v)\log P(v)$ where $P(v)=\sum_{x\in IS,dec(x)=v}P(x)$. $H_{ntp}$ is a coarse-grained Entropy compared to $H_{IS}$. $H_{IS}$ contains important information of previous tokens that are important for future tokens prediction, while $H_{ntp}$ is only related to next token.
Experiments in Figure 11 shows that, no matter what subspace we use, the Cross Entropy Loss usually follows a linear relationship with the Entropy we measured in the subspace, supporting the claim that the Next-Token-Prediction task likely lies in some subspace of the Intrinsic Space, or (statistically) its Entropy should be some weighted average of Entropy of several subspaces of similar dimension.. This also suggests $H_{ntp}$ is approximately linear with $H_{IS}$, thus validating our previous assumptions and claims.
Figure 11: Upper-left, Upper-right, Bottom-left: Cross Entropy Loss v.s. $\sum_{i\leq N}\log$ rel_eig_val; Bottom-right: correlation between minus CE loss and $\sum_{i\leq N}\log$ rel_eig_val. All experiments are for LLaMa-3.1-8B on a subset of OpenWebText. From the first three figure, we see CE loss is linear with the Entropy of certain subspaces. From the bottom-right figure, we see that Entropy measured in different subspaces are highly correlated ($corr>0.97$), which are also highly correlated with the CE loss for Next Token Prediction.
@@ C.2 Bridging the gap between Intrinsic Dimension explanation and Entropy explanation
Here, starting from previous assumptions and measurements w.r.t. Entropy in Intrinsic Space, we explain why CE is linear w.r.t. Intrinsic Dimension measured in Section 2.2, for $idx>500$. We see from Figure 3 that for $idx>500$, the eigen values mainly follow an exponential decay:
@@
We also now from previous result that for different context lengths, the relative eigen values increases almost at the same proportion, especially for $idx>1000$. That is, $\alpha_{ctl}=\alpha$, and $releigval_{ctl,\ idx}=\gamma(ctl)*releigval_{\infty,idx}$.
For the subspace for Next Token Prediction task, the Entropy should be proportional to log of volume in the subspace, hence it should be proportional to $m*\log\gamma(ctl)$ where $m$ is the dimension of this exact subspace. That is:
@@
For some certain threshold $thres$ we use, for some context length, the measured intrinsic dimension is:
@@
Thus deriving our assumptions in Section 2, where $s=m\alpha$. Equation (13) can also be validated in the fourth part of Figure 4, where the measured Intrinsic Dimensions (for $idx\geq 500$) are measured in the exponential-decaying area, and they share similar slope w.r.t. CE Loss.
@@ C.4 New Synthetic dataset: further experiments for Entropy
In the main paper we use $T<ctl$, where the task number is smaller than the context bits. Here, we further study the case where $T>ctl$. Moreover, instead of letting tasks show at the same frequency, we use weighted frequency proportional to $max(bit_{1},bit_{2})^{-1.2}$ as the weights of tasks.
@@
During training, $50\%$ of the samples are unmasked, while for the other $50\%$ samples we mask the last $X$ task bits to be $0.5$, where $X$ is a random int from $10$ to $59$. This ensures our model to be able to handle mask bits, and also ensures it can learn uncommon tasks (relying on context bits that are at the end of the context bits) well.
The context bits MLP has hidden size $400$, and the prediction MLP has hidden size $300$. The linear embedding and the context feature has same dimension $200$, and they are added before sending into the predict MLP. LeakyRELU with leaky hyper-parameter $31e-$ 3 is used as activation function for both MLPs. We train the model on a training set of $10000000$ and a validation set of size $1000000$, for $125$ epochs (and an early stopping setting of $25$ epochs, though the training process did not trigger early stopping).
To make sure that the model trained can be used to approximate the Bayes Model, we compare the model’s loss on validation set with context $ctl$ with the calculated minimum possible CE Loss for the task:
@@
And we obtain such result:
@@
Figure 13: Eigen value and CE results measured on trained model for Synthetic Dataset in this section. Left: relative eigen value v.s. Index. Middle: Model CE Loss v.s. $\sum_{i}\log rel\_eigval_{i}$ and Theoretical Min CE Loss v.s. $\sum_{i}\log rel\_eigval_{i}$. Right: Correlation between minus Experimentally measured model CE Loss, minus Theoretically Minimum CE Loss, and $\sum_{i\leq N}\log rel\_eigval_{i}$ for different $N$s. All correlation factors are larger than $0.97$.
@@ Appendix D More experiments of LLaMa on another dataset
Figure 14: Cross Entropy Loss v.s. Context Length, with log scale. We see that $y=C_{0}+C/x^{\gamma}$ fits this curve well.
According to Figure 14, we see $CE=C_{0}+C/l^{\gamma}$ approximates well for text corpora that is sured not to be seen by the model.

### v2

@@ 4.5 Optimal Context Length and Training Dataset Size: Synthetic Dataset
To validate the idea that longer context lengths are suitable for larger datasets, we conduct experiments on our proposed synthetic dataset with a fixed size MLP. During our experiments, we change the dataset size and context length used, and obtain validation Binary Cross Entropy loss vs. Context Length, as shown in the lower figure in Figure 1. (We also plot $y=-\log(1/2)$ as the loss of random guess in Figure 1.)
From Figure 1, we see that models with longer context length need larger datasets to train, and training long-context models on small datasets could lead to worse performance or even failure to improve from the first step (performance equal to random guess). Moreover, for each curve that represents a certain size of the training dataset used, we observe an optimal horizon beyond which validation loss would increase. This critical point moves right (optimal context length increases) with the increase of the training dataset size. This proves point 3 in Section 4.1.
@@ 5.1 Enlarging Context Length for LMs
Other work modifying architectures has also been proposed to enhance long context modeling or to simplify deployment of long context LLMs. For example, (Tang et al., 2024) proposes a training-free RazorAttention algorithm to largely compress the KV cache while maintaining performance unchanged.
Architectures and inference methods have been proposed to reduce inference time and memory cost for Language Models, represented by a series of linear transformer or RNN-based methods.(Katharopoulos et al., 2020; Gu & Dao, 2024; Sun et al., 2024). These methods, largely reducing the computational cost and memory usage for long input contexts, have displayed margin ahead of traditional attention-based langauge models for long context inference.
Currently a common practice to train very large Language Models supporting long context is to use pretrain the model with shorter contexts, then finetune them with longer contexts, as presented in tech reports of LLaMa-3(Grattafiori et al., 2024) and DeepSeek-v3(DeepSeek-AI et al., 2024).
@@ 5.2 Irrelevant Long Context hurts performance of LMs
Besides context length scaling with relevant contexts, previous researches have studied how LLMs perform for long irrelevant contexts. As an example, (Levy et al., 2024) studies the performance of current LLMs on an adjusted version of ‘needle in a haystack task, where two pieces of key information are embedded into a long text corpora and a question related to both is asked, similar to that presented in Figure 6. The conclusion of these work is that LLMs would perform worse when there is too much irrelevant information.
@@ 5.3 Long Context in another field: Time Series Forecasting
Context length, representing the length of input context, is not unique to Nature Language. For time series forecasting, where machine learning plays an important row, there is also work discussing the impact of context length, represented by (Shi et al., 2024). These investigations find that there exists an optimal look-back horizon, which increases with dataset size. However, time series datasets are relatively small compared to NLP datasets, and thus whether this conclusion holds on NLP remains an open problem for this work to study.
@@ 5.4 Related Theories for Scaling Laws
Since the discovery of Scaling Laws for Large Language Models (Kaplan et al., 2020) or even earlier, there has been theoretical work trying to explain why model performance could benefit from more data points and more model parameters. For exmaple, (Sharma & Kaplan, 2022) studies the dataset and model scaling from the data manifold perspective.
Specially for Language Models, there is also previous work proposing all kinds of theoretical models. For example, (Michaud et al., 2024) proposes a feature-quant based theory; (Aghajanyan et al., 2020) views the effect of fine-tuning from the intrinsic dimension perspective; (Havrilla & Liao, 2024) proposes to understand scaling with intrinsic dimensions.
@@ 6.1 Conclusion
In this work, we discuss the impact of context length on language modeling, especially Bayes risk and approximate loss, from both a theoretical and experimental perspective.
In Section 2, we propose assumptions related to the relationship between CE Loss, intrinsic dimension and context length. We derive a linear relation between CE loss and Intrinsic Dimension, and study the impact of context length to intrinsic dimension. We further investigate the relationship between intrinsic dimension, context length, and Entropy in intrinsic space in the appendix C from an Entropy perspective.
@@ 6.2 Limitations and Future Work
In Section 3 we mainly discuss the case for pretraining with long context, but common practice is to train with shorter contexts and then do long-context-training with longer contexts (DeepSeek-AI et al., 2024). We leave these to future work.
Our theory starting from Intrinsic Dimension only holds with assumptions in Section 2; and in Appendix C we use the perspective of Entropy and possible states number in Intrinsic Space to (partially) explain our assumptions and measurements w.r.t. Intrinsic Dimension. We hope future work may try to propose even more fundamental theories to explain our Intrinsic Dimension measurements and assumptions.
@@ Appendix A Definition of Cross-Entropy loss used in this work
It is well-known that
@@
This is because we assume that each dimension in the intrinsic space can store $S$ bits of information.
Another intuitive example is: assuming that the vocab is an integer from 00 to $12^{dim(\infty)*S}-$ 1. assuming $P(x_{0}|x_{-\infty:0})=\delta_{x_{0},y}$, that is, the next token given $x_{-\infty:0}$ is sure to be $y$. $y$. For $P_{l}(x_{0}|x_{-\infty:0})$, the first $dim(l)S$ digits of the integer (in binary representation) are known, but the remaining $(dim(\infty)-dim(l))S$ digits are unknown, making a guess in these numbers yield $P_{l}(x_{0}|x_{-\infty:0})=1/2^{S*(dim(\infty)-dim(l))}$. Thus, $D_{KL,x_{0}}(P(x_{0}|x_{-\infty,0}),P_{l}(x_{0}|x_{-\infty,0}))=1*\log 1/(1/2^{S*(dim(\infty)-dim(l))})=S*(dim(\infty)-dim(l))$.
@@ C.1 Bayes Risk from an Entropy-in-Intrinsic-Space perspective
Here we derive similar results as in Section 2.2, but not from an Intrinsic-Dimension perspective, rather we derive it from an Entropy perspective.
@@
  • Assumption 2. $\forall l_{1},l_{2}\text{ such that }l_{1}<l_{2}$, $H_{ntp}(l_{1})<H_{ntp}(l_{2})$. This is because a longer context contains more information about the next possible token.
@@
  • Assumption 3. ${D_{KL}}_{,ntp}(P,P_{l})\approx H_{ntp}(P)-H_{ntp}(P_{l})+Const$. This means that the KL divergence between the Bayes LM of infinite context length and the Bayes LM of context length $l$ can be written approximately in the form of entropy difference.
@@
To establish a relationship between Cross Entropy and Intrinsic Space, we run LLaMa-3-8b on a subset of the Openwebtext dataset and obtain the feature of the last token as the feature representation, or Intrinsic Space of the approximated Bayes Model, as shown in the left part of Figure 10. We see that the model with larger context length tends to have larger relative eigenvalues in intrinsic space, thus containing more information.
@@
Here $h$ is the ‘plank constant’, meaning that one state corresponds to a unit hyper-volume of $h^{dim}$ in the Intrinsic Space. A different value of $h$ would only add a constant to $H_{IS}$ and would not affect change in Entropy. Thus, we use $\sum_{i}\log rel\_eigval_{i}$ as Entropy in Intrinsic Space. Moreover, experiments in Figure 10 show that relative eigenvalues increase in the same ratio as the context length increases. This means the number of states in any subspace would scale proportionally with respect to the number of states in the whole space, that is: $H_{subspace}=H_{IS}*dim(subspace)/dim(IS)$, hence Entropy of subspaces are linear with respect to each other.
Though related, Entropy in Intrinsic Space does not equal to Entropy in the next token prediction task. From the probability perspective, let $dec(x)$ be the next decoded token for some point $x$ in the intrinsic space, we have: $H_{IS}=\sum_{x\in IS}-P(x)\log P(x)$, while $H_{ntp}=-\sum_{v\in vocab}P(v)\log P(v)$ where $P(v)=\sum_{x\in IS,dec(x)=v}P(x)$. $H_{ntp}$ is a coarse-grained Entropy compared to $H_{IS}$. $H_{IS}$ contains important information on previous tokens that are important for the prediction of future tokens, while $H_{ntp}$ is related only to the next token.
Experiments in Figure 11 show that, no matter what subspace we use, the Cross Entropy Loss usually follows a linear relationship with the Entropy we measured in the subspace, supporting the claim that the next token prediction task likely lies in some subspace of the Intrinsic Space, or (statistically) its Entropy should be some weighted average of Entropy of several subspaces of similar dimension.. This also suggests that $H_{ntp}$ is approximately linear with $H_{IS}$, which validates our previous assumptions and claims.
Figure 11: Upper-left, Upper-right, Bottom-left: Cross Entropy Loss vs. $\sum_{i\leq N}\log$ rel_eig_val; Bottom-right: correlation between minus CE loss and $\sum_{i\leq N}\log$ rel_eig_val. All experiments are for LLaMa-3.1-8B on a subset of OpenWebText. From the first three figure, we see CE loss is linear with the Entropy of certain subspaces. From the bottom-right figure, we see that Entropy measured in different subspaces are highly correlated ($corr>0.97$), which are also highly correlated with the CE loss for Next Token Prediction.
@@ C.2 Bridging the gap between Intrinsic Dimension explanation and Entropy explanation
Here, starting from previous assumptions and measurements w.r.t. Entropy in Intrinsic Space, we explain why CE is linear w.r.t. Intrinsic Dimension measured in Section 2.2, for $idx>500$. We see in Figure 3 that for $idx>500$, the eigenvalues mainly follow an exponential decay:
@@
We also now from the previous result that for different context lengths, the relative eigenvalues increase almost in the same proportion, especially for $idx>1000$. That is, $\alpha_{ctl}=\alpha$, and $releigval_{ctl,\ idx}=\gamma(ctl)*releigval_{\infty,idx}$.
For the subspace for the next token prediction task, the entropy should be proportional to log of volume in the subspace; hence it should be proportional to $m*\log\gamma(ctl)$ where $m$ is the dimension of this exact subspace. That is:
@@
For some certain threshold $thres$, we use, for some context length, the measured intrinsic dimension is:
@@
Thus, we derive our assumptions in Section 2, where $s=m\alpha$. Equation (13) can also be validated in the fourth part of Figure 4, where the measured Intrinsic Dimensions (for $idx\geq 500$) are measured in the exponential decay area, and share similar slopes w.r.t. CE Loss.
@@ C.4 New Synthetic dataset: further experiments for Entropy
In the main paper, we use $T<ctl$, where the task number is smaller than the context bits. Here, we further study the case where $T>ctl$. Moreover, instead of letting tasks show at the same frequency, we use weighted frequency proportional to $max(bit_{1},bit_{2})^{-1.2}$ as the weights of tasks.
@@
During training, $50\%$ of the samples are unmasked, while for the other $50\%$ samples, we mask the last $X$ task bits to be $0.5$, where $X$ is a random int from $60-10$ to $60-60$. This ensures our model to be able to handle mask bits, and also ensures it can learn uncommon tasks (relying on context bits that are at the end of the context bits) well.
The context bits MLP has hidden size $400$, and the prediction MLP has hidden size $300$. The linear embedding and the context feature have the same dimension $200$, and are added before being sent to the predict MLP. LeakyRELU with leaky hyperparameter $31e-$ 3 is used as an activation function for both MLPs. We train the model on a training set of $10000000$ and a validation set of size $1000000$, for $125$ epochs (and an early stopping setting of $25$ epochs, though the training process did not trigger early stopping).
To make sure that the trained model can be used to approximate the Bayes Model, we compare the model’s loss on validation set with context $ctl$ with the calculated minimum possible CE Loss for the task:
@@
And we obtain such a result:
@@
Figure 13: Eigen value and CE results measured on trained model for Synthetic Dataset in this section. Left: relative eigen value vs. Index. Middle: Model CE Loss vs. $\sum_{i}\log rel\_eigval_{i}$ and Theoretical Min CE Loss vs. $\sum_{i}\log rel\_eigval_{i}$. Right: Correlation between minus Experimentally measured model CE Loss, minus Theoretically Minimum CE Loss, and $\sum_{i\leq N}\log rel\_eigval_{i}$ for different $N$s. All correlation factors are larger than $0.97$.
@@ Appendix D More experiments of LLaMa on another dataset
Figure 14: Cross Entropy Loss vs. Context Length, with log scale. We see that $y=C_{0}+C/x^{\gamma}$ fits this curve well.
According to Figure 14, we see that $CE=C_{0}+C/l^{\gamma}$ approximates well for text corpora that are sure not to be seen by the model.